OpenStudy (anonymous):

find the solution to the system of equations by the substitution method. -5x-7y=-57 3x-2y=-3

OpenStudy (precal):

3x-2y=-3 solve for x x=-1+2/3y sub into the first equation solve for y

OpenStudy (anonymous):

i don't get it :/

hero (hero):

-5x-7y=-57 3x-2y=-3 Isolate y for each equation: y = (5x - 57)/-7 y = (3 + 3x)/2 Set y =y, solve for x: 2(5x - 57) = -7(3 + 3x) 10x - 114 = -21 -21x 10x + 21x = 114 - 21 31x = 93 x = 93/31 x = 3 y = 6

hero (hero):

@brittwhit18

OpenStudy (precal):

Hero's approach will also, work. The whole idea of this method is to substitute x= something into the other equation or substitue y=something into the other equation Many ways to accomplish the goal - same solution in the end